Question 4 1 pts A study is done to see if the average age a child moves permanently out of his parents home in the United
Answer #1

The correct option is: B 0.1738 The number of US adults is: n=50 The sample average age is: 1 = 26.1 The sample standard devi

Calculate the degrees of freedom. df =n-1 = 50-1 = 49 Calculate the p-value. p-value =(=TDIST(0.9485,49,1)) = 0.1738 (Use MS